A 3.00-N ball collides with a 10.0-m high wall. Calculate the impulse and size of the force of the wall for each of the following cases:

a. The ball’s speed is 5 m/s just before it hits the wall and 5 m/s just after leaving the wall. The time of collision is 0.100 s.

b. The ball’s speed is 5 m/s just before it hits the wall and 4 m/s just after leaving the wall. The time of collision is 0.100 s.

c. The ball’s speed is 5 m/s just before it hits the wall and 4 m/s just after leaving the wall. The time of collision is 0.500 s.

d. The ball’s speed is 5 m/s just before it hits the wall and it sticks to the wall (ball and wall are made of Velcro). The time of collision is 0.100 s.
Answer #1

I = FΔt = ΔP = m* ( vf -vi )

F = I/Δt

m = 3/9.8 kg

A)

I =  3/9.8*(5-(-5))   =3.06122 kg *m/s

F = 3.06122 / 0.100 =30.6122 N

B)

I =  3/9.8*(5-(-4))   =2.755 kg *m/s

F = 2.755 / 0.100 =27.55 ​​​​​​​ N

C)

I =  3/9.8*(5-(-4))   =2.755 kg *m/s

F = 2.755 / 0.500 =5.51 ​​​​​​​ N

D)

I =  3/9.8*(5-(-0))   =1.506 kg *m/s

F = 1.506 / 0.100 =15.06 ​​​​​​​ N

Goodluck for exams Comment in case any doubt, will reply for sure..